# Counting Distinct Elements in the Turnstile Model with Differential Privacy under Continual Observation

### v1

@@ 5.2 From the reduction to the accuracy lower bound
In different parameter regimes, we get lower bounds $\Omega(T)$ and $\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon\log w},\frac{T^{1/3}}{\varepsilon^{2/3}\log\varepsilon T}\right)\right)$. A similar proof works for $\delta=0$. Hence, overall, we get a lower bound of $\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon\log w},\frac{T^{1/3}}{\varepsilon^{2/3}\log\varepsilon T},T\right)\right)$. ∎
Appendix A A concentration inequality for Gaussian random variables
Lemma A.1.
@@ Appendix A A concentration inequality for Gaussian random variables
$\Pr[|R|>\ell]\leq 2e^{-\frac{\ell^{2}}{2\sigma^{2}}}.$
Lemma A.2.
@@
$\Pr[\max_{j\in[m]}|R_{j}|>\ell]\leq 2me^{-\frac{\ell^{2}}{2\sigma^{2}}}.$
@@
By a union bound and Lemma A.1,
$\Pr[\max_{i\in[m]}|R_{i}|>\ell]=\Pr(\exists i\in[m]\text{ such that }|R_{i}|>\ell)$
$\leq\sum_{i=1}^{m}\Pr(|R_{i}|>\ell)\leq\sum_{i=1}^{m}2e^{-\frac{\ell^{2}}{2\sigma^{2}}}=2me^{-\frac{\ell^{2}}{2\sigma^{2}}}.$
Appendix B General lower bound reduction for small $\varepsilon$
We describe an adaptation of a folklore reduction to our problem of interest that allows us to extend a lower bound for $\varepsilon=1$ to any $\varepsilon<1$. The theorem is stated for item-level differential privacy, but applies to event-level differential privacy as well.
Theorem B.1.
Let $\alpha:\mathbb{R}\to\mathbb{R}$ be an increasing function. Let $T\in\mathbb{N}$, and $\varepsilon,\delta\in[0,1]$ such that $\varepsilon\geq\frac{1}{T}$. If for all $T\in\mathbb{N}$, any mechanism for $\mathsf{CountDistinct}{}$ that achieves $(1,\delta)$-item-level DP for streams of length $T$ has error parameter at least $\alpha(T)$, then any mechanism for $\mathsf{CountDistinct}{}$ that achieves $(\varepsilon,\delta\varepsilon/10)$-item-level DP for streams of length $T^{\prime}$ has error parameter at least $\frac{\alpha(\varepsilon T^{\prime})}{2\varepsilon}$.
Proof.
Let $T^{\prime}\in\mathbb{N}$, $\varepsilon,\delta\in[0,1]$. Let $\ell=\lfloor 1/\varepsilon\rfloor$. We prove the contrapositive. Namely, let $\mathcal{M}^{\prime}$ be a mechanism for $\mathsf{CountDistinct}$ that is $\frac{\alpha(\varepsilon T^{\prime})}{2\varepsilon}$-accurate and $(\varepsilon,\frac{\delta\varepsilon}{10})$-item-level-DP for streams of length $T^{\prime}$. We will use $\mathcal{M}^{\prime}$ to construct a mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$ that is $\alpha(T)$-accurate and $(1,\delta)$-item-level-DP for streams of length $T=\lceil T^{\prime}/\ell\rceil$.
Given a stream $x\in\mathcal{U}_{\pm}^{T}$ of length $T=\lceil T^{\prime}/\ell\rceil$, we create a new stream $x^{\prime}$ of length $T^{\prime}$ with insertions and deletions from a larger universe $\mathcal{U}^{\prime}=[\ell]\times\mathcal{U}$ (so every item in $\mathcal{U}$ corresponds to $\ell$ distinct items in $\mathcal{U}^{\prime}$), as follows: Initialize $x^{\prime}$ to be empty. For $t\in[T]$, if $x[t]=+i$, append $+(i,1),\ldots,+(i,\ell)$ to $x^{\prime}$. Similarly, if $x[t]=-i$, append $-(i,1),\ldots,-(i,\ell)$. If $x[t]=\perp$ then, append $\ell$ stream entries ,$\perp$,$\dots$,$\perp$ , … , ⟂ to $x^{\prime}$. Finally, define the output of $\mathcal{M}$ as follows: run $\mathcal{M}^{\prime}$ on $x^{\prime}$ and, for each time step $t\in T$, output $\mathcal{M}(x)[t]=\frac{1}{\ell}\cdot\mathcal{M}^{\prime}(x^{\prime})[t\cdot\ell]$.
Changing a single item of $x$ changes at most $\ell$ items of $x^{\prime}$. Therefore, by the privacy guarantee of $\mathcal{M}^{\prime}$, and by group privacy, $\mathcal{M}$ achieves $(1,\delta)$-item-level DP, for streams of length $T=\lceil T^{\prime}/\ell\rceil$. Finally, the accuracy guarantee of $\mathcal{M}$ will follow from the accuracy guarantee of $\mathcal{M}^{\prime}$ : By the construction of $x^{\prime}$, it holds $\mathsf{CountDistinct}(x^{\prime})[t\cdot\ell]=\ell\cdot\mathsf{CountDistinct}(x)[t]$, and by the accuracy of $\mathcal{M}^{\prime}$, with probability $0.99$, the maximum error of $\mathcal{M}^{\prime}$ on any input stream $x^{\prime}$ of length $T^{\prime}$ is at most $\alpha(\varepsilon T^{\prime})/2\varepsilon$. Therefore, with probability at least $0.99$, the error of $\mathcal{M}$ is at most $\frac{1}{\ell}\cdot\alpha(\varepsilon T^{\prime})/2\varepsilon\leq\frac{1}{2\varepsilon\ell}\cdot\alpha(\varepsilon\cdot\ell\cdot T)\leq\alpha(T)$, as desired. This concludes the proof of Theorem B.1.  ∎

### v3

@@ 5.2 From the reduction to the accuracy lower bound
In different parameter regimes, we get lower bounds $\Omega(T)$ and $\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon\log w},\frac{T^{1/3}}{\varepsilon^{2/3}\log\varepsilon T}\right)\right)$. Overall, we get a lower bound of $\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon\log w},\frac{T^{1/3}}{\varepsilon^{2/3}\log\varepsilon T},T\right)\right)$.
Finally, consider $\delta=0$. Lemma 5.5 provides a lower bound of $\alpha^{\prime}=\Omega\Big{(}\mathsf{min}\left(w,\sqrt{T}\right)\Big{)}$ on accuracy for ($\varepsilon^{\prime}=1$, $\delta=0$)-item-level-DP, $\alpha^{\prime}$-accurate mechanisms for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$. By invoking Theorem A.1 with $\delta=0$, we can extend this to the following lower bound for $(\varepsilon,0)$-DP mechanisms:
$\alpha=\Omega\Big{(}\mathsf{min}\Big{(}\frac{w}{\varepsilon},\frac{\sqrt{T}}{\varepsilon}\Big{)}\Big{)}.$
Combining this bound with the $\Omega(T)$ lower bound for when $\varepsilon<\frac{2}{T}$ gives the desired result. ∎
Acknowledgments
We thank Teresa Steiner and an anonymous reviewer for useful suggestions on the initial version of this paper. S.S. was supported by NSF award CNS-2046425 and Cooperative Agreement CB20ADR0160001 with the Census Bureau. A.S. and P.J. were supported in part by NSF awards CCF-1763786 and CNS-2120667 as well as Faculty Awards from Google and Apple.
Appendix A General lower bound reduction for small $\varepsilon$
We describe an adaptation of a folklore reduction to our problem of interest that allows us to extend a lower bound for $\varepsilon=1$ to any $\varepsilon<1$. The theorem is stated for item-level differential privacy, but applies to event-level differential privacy as well. In addition, this reduction also works for variants of the model previously discussed: offline mechanisms, the likes model, and the strict turnstile model.
Theorem A.1.
Let $\alpha:\mathbb{R}\to\mathbb{R}$ be a nondecreasing function. Let $T\in\mathbb{N}$, and $\varepsilon,\delta\in[0,1]$ such that $\varepsilon\geq\frac{1}{T}$. If for all $T\in\mathbb{N}$, every mechanism for $\mathsf{CountDistinct}{}$ that is $(1,\delta)$-item-level-DP for streams of length $T$ has error parameter at least $\alpha(T)$, then every mechanism for $\mathsf{CountDistinct}{}$ that is $(\varepsilon,\frac{\delta\varepsilon}{2})$-item-level-DP for streams of length $T^{\prime}$ has error parameter at least $\frac{\alpha(\varepsilon T^{\prime})}{2\varepsilon}$.
Proof.
Fix $\varepsilon\in(0,1]$ and $\delta\in[0,1)$. Let $\ell=\lfloor 1/\varepsilon\rfloor$ and $T^{\prime}\in\mathbb{N}$, where $T^{\prime}$ is divisible by $\ell$. We prove the contrapositive. Namely, let $\mathcal{M}^{\prime}$ be a mechanism for $\mathsf{CountDistinct}$ that is $\frac{\alpha(\varepsilon T^{\prime})}{2\varepsilon}$-accurate and $(\varepsilon,\frac{\delta\varepsilon}{2})$-item-level-DP for streams of length $T^{\prime}$. We use $\mathcal{M}^{\prime}$ to construct a mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$ that is $\alpha(T)$-accurate and $(1,\delta)$-item-level-DP for streams of length $T={T^{\prime}/\ell}$.
Given a stream $x\in\mathcal{U}_{\pm}^{T}$ of length $T={T^{\prime}/\ell}$, create a new stream $x^{\prime}$ of length $T^{\prime}$ with insertions and deletions from a larger universe $\mathcal{U}^{\prime}=\mathcal{U}\times[\ell]$ (so every item in $\mathcal{U}$ corresponds to $\ell$ distinct items in $\mathcal{U}^{\prime}$), as follows: Initialize $x^{\prime}$ to be empty. For $t\in[T]$, if $x[t]=+i$, append $+(i,1),\ldots,+(i,\ell)$ to $x^{\prime}$. Similarly, if $x[t]=-i$, append $-(i,1),\ldots,-(i,\ell)$. If $x[t]=\perp$, append $\ell$ stream entries $\perp$ to $x^{\prime}$. Finally, define the output of $\mathcal{M}$ as follows: run $\mathcal{M}^{\prime}$ on $x^{\prime}$ and, for each time step $t\in T$, output $\mathcal{M}(x)[t]=\frac{1}{\ell}\cdot\mathcal{M}^{\prime}(x^{\prime})[t\cdot\ell]$.
Changing a single item of $x$ changes $\ell$ items of $x^{\prime}$. Therefore, by the privacy guarantee of $\mathcal{M}^{\prime}$, by group privacy (see Lemma 2.3) and since $e^{\varepsilon}-1\geq\varepsilon$ for all $\varepsilon$, we get that $\mathcal{M}$ is $(\varepsilon^{*},\delta^{*})$-item-level-DP for streams of length $T={T^{\prime}/\ell}$, where $\varepsilon^{*}=\ell\varepsilon\leq 1$ and $\delta^{*}=\frac{\delta\varepsilon}{2}\cdot\frac{e^{\ell\varepsilon}-1}{e^{\varepsilon}-1}\leq\frac{\delta\varepsilon}{2}\cdot\frac{2}{\varepsilon}=\delta$. Thus, $\mathcal{M}$ is $(1,\delta)$-item-level-DP for streams of length $T$.
Finally, we derive the accuracy guarantee of $\mathcal{M}$ from the accuracy guarantee of $\mathcal{M}^{\prime}$. By the construction of $x^{\prime},$
$\mathsf{CountDistinct}(x^{\prime})[t\cdot\ell]=\ell\cdot\mathsf{CountDistinct}(x)[t].$
By the accuracy of $\mathcal{M}^{\prime}$, on every input stream $x^{\prime}$ of length $T^{\prime}$, with probability at least $0.99$, the maximum error of $\mathcal{M}^{\prime}$ is at most $\frac{\alpha(\varepsilon T^{\prime})}{2\varepsilon}$. Therefore, with probability at least $0.99$, the maximum error of $\mathcal{M}$ is at most $\frac{1}{\ell}\cdot\frac{\alpha(\varepsilon T^{\prime})}{2\varepsilon}\leq\frac{1}{2\varepsilon\ell}\cdot\alpha(\varepsilon\cdot\ell\cdot T)\leq\alpha(T)$, as desired. This concludes the proof of Theorem A.1. ∎
Appendix B Lower bounds for variants of the model
In this section, we prove lower bounds for $\mathsf{CountDistinct}$ in more restricted turnstile streaming models. In Theorem B.3, we prove that our item-level lower bound applies to the more restricted likes model, whereas in Theorem B.5, we prove that both our event-level and item-level lower bounds apply to the strict turnstile model. In addition, our lower bounds apply even when considering only offline mechanisms for $\mathsf{CountDistinct}$, defined next.
Definition B.1 (Offline mechanisms for $\mathsf{CountDistinct}$).
A mechanism for $\mathsf{CountDistinct}$ is an offline mechanism if it first receives the entire input stream $x$ before outputting the vector of $\mathsf{CountDistinct}$ estimates.
Definition B.2 (Likes model).
In the likes model, an element can be inserted into the stream only when it is absent and deleted from the stream $x$ only when it is present. Mathematically, for a stream $x$ of length $T$, for all $t\in[T]$, if $x[t]=+u$ for some element $u\in\mathcal{U}$, then $f_{u}(x)[t-1]=0$, and if $x[t]=-u$, then $f_{u}(x)[t-1]=1$, where we define $f_{u}(x)[0]=0$ for convenience.
Theorem B.3 (Item-level lower bound in the likes model).
The item-level lower bound in Theorem 1.7 holds for all offline mechanisms for $\mathsf{CountDistinct}$ in the likes model.
Proof.
The proof proceeds identically to the proof of Theorem 1.7, with the following observations. First, note that our reduction in Algorithm 5 uses the answers returned by the mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$ only after $\mathcal{M}$ has produced the entire stream of answers (it does not require online access to the outputs of $\mathcal{M}$). Therefore, one can restate the reduction so that it only requires black-box access to an offline mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$, and the proof of Theorem 1.7 would still hold.
Next, observe that the stream $x$ constructed in the reduction in Algorithm 5 belongs to the likes model. This holds because, for each element, the stream entries belonging to that element alternate between insertions and deletions, with the first entry always being an insertion. Finally, the lower bound reduction from $\varepsilon=1$ to $\varepsilon<1$ in Theorem A.1 works both for offline mechanisms and streams in the likes model. ∎
Definition B.4 (Strict turnstile model).
In the strict turnstile model, an element can be deleted only when it is present in the stream. Mathematically, for a stream $x$ of length $T$, for all $t\in[T]$, if $x[t]=-u$ for some time step $t$ and element $u\in\mathcal{U}$, then $f_{u}(x)[t-1]=1$, where we define $f_{u}(x)[0]=0$ for convenience.
Theorem B.5 (Lower bounds for the strict turnstile model).
The event-level lower bound of Theorem 1.6 and the item-level lower bound of Theorem 1.7 hold for all offline mechanisms for $\mathsf{CountDistinct}$ in the strict turnstile model.
Proof.
Since the likes model is a special case of the strict turnstile model, by Theorem B.3, the item-level lower bound applies to the strict turnstile model as well. We now focus on the event-level lower bound. Again, the proof proceeds identically to the proof of Theorem 1.6, with the following observations.
First, note that our reduction in Algorithm 4, similarly to the reduction in Algorithm 5, uses the outputs of the mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$ only after the mechanism has produced the entire stream of answers. Therefore, one can restate the reduction so that it only requires black-box access to an offline mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$.
Next, the stream $x$ constructed in Step 10 of the reduction in Algorithm 4 belongs to the strict turnstile model, since each element deletion is preceded by an insertion of the same element. Finally, the lower bound reduction from $\varepsilon=1$ to $\varepsilon<1$ in Theorem A.1 works both for offline mechanisms and streams in the strict turnstile model. ∎
Appendix C A concentration inequality for Gaussian random variables
Recall that $\mathcal{N}(\mu,\sigma^{2})$ denotes the Gaussian distribution with mean $\mu$ and standard deviation $\sigma$.
Lemma C.1.
@@ Appendix C A concentration inequality for Gaussian random variables
$\Pr[|R|>\lambda]\leq 2e^{-\frac{\lambda^{2}}{2\sigma^{2}}}.$
Lemma C.2.
@@
$\Pr[\max_{j\in[m]}|R_{j}|>\lambda]\leq 2me^{-\frac{\lambda^{2}}{2\sigma^{2}}}.$
@@
By a union bound and Lemma C.1,
$\Pr[\max_{j\in[m]}|R_{i}|>\lambda]$ $=\Pr[\exists i\in[m]\text{ such that }|R_{i}|>\lambda]$
[…]